OpenStudy (abbycross167):

Can someone please show me how to answer this question? Please? Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. 9x + 12y= -36

OpenStudy (mathmale):

Your choice: point-slope method or intercepts method?

OpenStudy (abbycross167):

@mathmale which do you think will be easier ?

OpenStudy (mathmale):

For 9x + 12y= -36 I'd use the intercept method. Let x = 0 and find y. Then the y-intercept is the point (0, ? ) Let y = 0 and find x. Then the x-intercept is the point ( ? , 0) Plot these on coordinate axes and draw a line thru both points.

OpenStudy (mathmale):

y-intercept: Let x = 0 and find y. Letting x=0, 9(0) + 12y = -36. What is y?

OpenStudy (abbycross167):

24? @mathmale

Directrix (directrix):

9(0) + 12y = -36 0 + 12y =- 36 Try again @abbycross167

OpenStudy (abbycross167):

-24? @Directrix

Directrix (directrix):

I think you are adding 12 to -36. That is not what the equation is asking you to do. 12y =- 36 Divide both sides by 12. What is -36 divided by 12?

OpenStudy (abbycross167):

-3 @Directrix

Directrix (directrix):

Yes. Let me read through the thread to see what we are supposed to do next.

Directrix (directrix):

For 9x + 12y= -36 I'd use the intercept method. Let x = 0 and find y. Then the y-intercept is the point (0, -3 ) Let y = 0 and find x. Then the x-intercept is the point ( -4 , 0) Plot these on coordinate axes and draw a line thru both points.

Directrix (directrix):

Click on the pencil and sketch the line in blue. Recall that two points determine a line.